OpenStudy (anonymous):

how do i graph y=-1 over x + 1 on a graph??

OpenStudy (anonymous):

you mean you want to sketch y= -1 on a graph right?

OpenStudy (anonymous):

im not sure, thats why im looking for help where ever i can get it. i have an assignment due wednesday and that is one of the questions. the teacher is very little help and hasnt told me what to do. can you help?

OpenStudy (anonymous):

what does the question really say?

OpenStudy (anonymous):

the question says: on graph D draw y=-1 over x+1

OpenStudy (anonymous):

graph D is x + 1 right?

OpenStudy (anonymous):

well i know it isnt equal to negative one.

OpenStudy (anonymous):

hmmm if so, then you have point y as -1 and you can find x : y = x + 1 -1 = x + 1 x = -2 so , your point on that graph must be at (-2,-1) I guess.

OpenStudy (anonymous):

since you want to graph y= -1 on x + 1, then they are equal at some point

myininaya (myininaya):

? when you over do you mean it is a fraction like -1/(x+1)

OpenStudy (anonymous):

I think that's what he meant! then I have misread the question again lol, sorry

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then @matto forget my answer, I have misread the question, dearest apology

myininaya (myininaya):

Do you know what 1/x looks like?

OpenStudy (anonymous):

hyperbola

OpenStudy (anonymous):

to graph such question you have to find the limits to find the HAs and VAs

myininaya (myininaya):

just move the graph over 1 in the positive direction

OpenStudy (anonymous):

yes, it is bacically a an open c shape that only takes up one quater of a quadratic graph. im not sure how t explain it

myininaya (myininaya):

sorry over 1 in the negative direction

OpenStudy (anonymous):

he can't do that without finding the HAs and VAs

OpenStudy (anonymous):

you have to find the limit : \[\lim_{x \rightarrow \pm \infty} y \] to find the HAs and \[\lim_{x \rightarrow -1\pm}\] to find the VAs

OpenStudy (anonymous):

to see where the function's position will be in the end

OpenStudy (anonymous):

that's it :)

OpenStudy (anonymous):

im sorry, but in all seriousness, i have absolutly no idea what that means.

OpenStudy (anonymous):

lol, alright did you take limits?

OpenStudy (anonymous):

thank you! actually I have been strugling with this problem too... thanks now i am starting to get it... btw, I think you could be a GREAT lecturer

OpenStudy (anonymous):

thanks andy ^_^ I'm flattered

OpenStudy (anonymous):

maths is easily my weakest subject, im doing 4 3A/B subjects for year 12 and i was hoping this migh help.

OpenStudy (anonymous):

don't worry, it will help. I was weak in mathematics , but then I got on my feet and worked hard and look where have I reached :)

OpenStudy (anonymous):

@ matto, did you take limits?

OpenStudy (anonymous):

what is a limit?

OpenStudy (anonymous):

this :\[\lim_{a \rightarrow b} f(x) \] = L do you recognize this?

OpenStudy (anonymous):

im sorry, no. as i have said, my teacher has done very little teaching over the last 8 weeks.

OpenStudy (anonymous):

oh, btw, I have a book called "calculus" =D its like 1000 pages, anyone want it? If it feels too big, I also have Calculus Essentials, maybe you want this?

OpenStudy (anonymous):

andy, can you explain the problem for him please? :) brb

OpenStudy (anonymous):

the problem looks simple, but also looks uncomplete... I cant really get what the question means..

OpenStudy (anonymous):

it also asks to include all asymptotes if tha helps. i dont know what an asymptote is.

OpenStudy (anonymous):

In analytic geometry, an asymptote of a curve is a line such that the distance between the curve and the line approaches zero as they tend to infinity.

OpenStudy (anonymous):

ok, well, its ok if you cant help. i appreciate the help. this was basically a last resort for me, i will seek help else where.

OpenStudy (anonymous):

you know, i think you should go to your friends ask them to help if they dont, go to your teacher if she doesnt, go to library

OpenStudy (anonymous):

if library doesnt, drop it, go do something fun...

OpenStudy (anonymous):

back, alright @matto

OpenStudy (anonymous):

i will ask my friends

OpenStudy (anonymous):

An asymptote is a line you draw on the graph that limits the graph from extending , contracting etc we have 2 types of asymptotes 1) Horizontal asymptotes are the lines that cut the y-axis horizontaly, you find them by taking the limit of f(x) as x goes to infinity like this: \[\lim_{x \rightarrow \pm \infty} f(x)\] and you take the limit from the left and right, if you end up with the limit = : if you end up with an answer that is a number, then that number is you Horizontal asymptote. 2) Vertical asymptotes are the lines that cut the x-axis, and you find them by taking the limit of x, as x goes to a number that makes your denominator 0, in your function which is y = -1/x+1 , x=-1 your number : \[\lim_{x \rightarrow 1\pm} f(x)\] \[\infty , -\infty\] then x = 1 is your Vertical Asymptote

OpenStudy (anonymous):

I hope he got the meaning and difference between HAs and VAs

OpenStudy (anonymous):

i hope so too

OpenStudy (anonymous):

it's not simple to explain it without showing him the graph by hand

OpenStudy (anonymous):

x = -1 * instead of x=1 sorry ^^"